OpenStudy (shelby1290):

find the slope of the line by writing in the form y=mx+b. x-2y=5

OpenStudy (anonymous):

2y=x-5

OpenStudy (anonymous):

y=x/2-5/2

OpenStudy (anonymous):

is in the form y=mx+b

OpenStudy (shelby1290):

@d3v13 so i just do 1 divided by 2 minus five divided by two

OpenStudy (shelby1290):

i dont quite get it ...sorry

OpenStudy (anonymous):

no. slope is m in the equation

OpenStudy (anonymous):

see what is equivalent to m in our equation

OpenStudy (anonymous):

y=mx +b y=(1/2)x+ (-5/2)

OpenStudy (shelby1290):

oh so you just substitute everything

OpenStudy (shelby1290):

and move the b to the other side

OpenStudy (shelby1290):

right?

OpenStudy (anonymous):

m=1/2= slope

OpenStudy (shelby1290):

okay

OpenStudy (anonymous):

sorry I was not trying to be rude. But I don't know how to explain further. my English is not that good

OpenStudy (shelby1290):

oh no, its fine.